OpenStudy (anonymous):

A gardener decides to enclose a rectangular garden. He has 32 feet of fence for this purpose. If his garden is x feet long, express the area of his garden as a function of the length.

OpenStudy (anonymous):

I have the answere just want to know how to do it

OpenStudy (anonymous):

the area should be the multiplication of length times width. Since he only has 32 feet, 32 is equal to the length of all the sides (2x+2W) where w is the width. That is the first equation. The second is the area equation Area=x*W. solve the first equation for W and insert in the second equation 32=2x+2W, divide by two, the subtract x from both sides. 32/2 - x = W. Put this back into the area equation and you have Area=x*(32/2 - x)